I am 27 years older tan my son. In 14 years time i will be twice his age. When were we born?

Answer by Cromlix(4381)

Hi there,
I am 27 years older than my son
I = 27 + S..........(1)
In 14 years time i will be twice his age
I + 14 = 2(S + 14)......(2)
Replace I in Equation (2) with
27 + S from Equation (1)
I + 14 = 2(S + 14)
27 + S + 14 = 2S + 28
Collect like terms:
S - 2S = 28 -27 - 14
-S = -13
S = 13
I am currently 40 years old.
My son is currently 13 years old.
Hope this helps :-)